OpenStudy (anonymous):

What is the equation of y=x^3 with the given transformations? vertical stretch by a factor of 3, horizontal shift 4 units to the right, vertical shift 3 units down

OpenStudy (anonymous):

do you know what a, b and c mean in this function: F(x)=a(x-b)^3+c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so what it will be after stretch?

OpenStudy (anonymous):

f(x)=3x^3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then f(x)=3*(x-4)^3

OpenStudy (anonymous):

and finally f(x)=3*(x-4)^3-3